OpenStudy (anonymous):

could someone please help me with this question? If 3x − 4 ≤ f(x) ≤ x^2 − 3x + 5 for x ≥ 0, find lim of f(x) as x→3

ganeshie8 (ganeshie8):

take limit through out \[\large \lim \limits_{x\to 3} 3x − 4 ≤ \lim \limits_{x\to 3} f(x) ≤ \lim \limits_{x\to 3} x^2 − 3x + 5 \]

ganeshie8 (ganeshie8):

evaluate the limit for left and right expressions ^^

OpenStudy (anonymous):

so limit = 5 ?

ganeshie8 (ganeshie8):

yes ! \[\large 5≤ \lim \limits_{x\to 3} f(x) ≤ 5\] gives you \(\large \lim \limits_{x\to 3} f(x) = 5 \)

OpenStudy (anonymous):

thanks alot